Problem 7.4. Solve the following IVPs. 2t 10 e 8.e a) 7.y10e Yo = 4 b) y- 2.8. y 6.e, -2, yo -1 8 y 6. e4t. yo c) y-5.4y 3
Answer #1

8 e InteytinFeeta L- 2t-it toJet 1e -St Ye MP 8 C - 8 -e ce Giver o) =4 -2 -1 + C C1 -t 2t - e t7e 2e)-2- Auille etion D- - 8 o P= -2, 4 Complementory Fundtion c ie Ce P2D-8 -6t e -te 2D-2 8-2 iP-e ut +te .4=Ce Gimen yCo)ry C